08 September 2015 UT (wake-up, revisited...)

OSA: David Gonzalez Huerta

ISp: John Morris

Telescope Manager: Jeff Urban

ISAs: Olga Kuhn, Barry Rothberg

Another try at on-sky checkout. Clouds and high humidity at sunset, but clearing and humidity dropping now (3 UT).

LBCs in beam and authorized.

03 UT: Series of 10 biases on LBCB and LBCR.

03:09 UT: Opening shutter doors. Jeff and Dan on standby. Closing again to check all is working.

03:30 UT: Opening shutter doors "for real".

03:35 UT: Copoint field for IC4665 - id'd field.

03:40 UT: dofpia, /x2,/first
  • hung on waiting for images in /newdata.
  • timestamp of LBCR exposures is way off! 20150903 and 18:34 vs 20150908 and 03:40 for LBCB.
    • called Kellee. We used vncviewer to open RSci PC. Right-click on time at bottom right. Go to Date&Time tab, manually set date to Sept 8 and time close to current UT. Go to Internet Time tab. Click Update Now. Sync'ed time on all 4 PCs, although Red Sci was the only one that looked obviously off.
  • ran dofpia again after syncing times (~04 UT). Not hanging any longer.

04:00 UT we closed for rising humidity and clouds.

07:25 UT - reopened and slewed to copoint star for M103 field.

07:35 UT - Copointing completed - LBC-B looks elongated, rerunning DOFPIA

07:39 UT - Running NoDither M103 script to check rotation center, PA=0,180, then PA=90,270. First script sends preset, reset have Dec=-90 PA=0 73952 (B), 73946(R) PA=180 74429 (B), 74424 (R) PA=90 74740 (B), 74734 (R) PA=270 75101 (B), 75055 (R)

07:52 UT - Refocus, DOFPIA

07:57 UT - Running Superfoc sequence for checking Tilt, RB_rv_SF7.ob for 4PAs (0,180 first)

08:04 UT - Running RB_rv_SF7_pa180.ob

08:12 UT - Running RB_rv_SF7_pa90.ob

08:19 UT - Running RB_rv_SF7_pa270.ob

08:28 UT - running DOFPIA

08:33 UT - Running M103 Dither sequence - M103_9ptL_PA0 (40" dither) - scaling exp 3x so every image is guiding -every filter, 45-180sec 9pt dither

08:42 UT - stars seem to be on a tea-break on the all-sky, hope they return soon

08:56 UT - Stopped exposure at 2/6(B) 3/7(R) - did one dither shift and ran through all the filters

08:57 UT - Closing up

11:27 UT - Turning off LBC and disconnecting from telescope

-- %USERSIG{OlgaKuhn - 2015-09-08}%

Comments

 
Topic revision: r3 - 08 Sep 2015, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback